Circular No. 5743

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

SUPERNOVA 1993J IN NGC 3031 R. A. Sramek, National Radio Astronomy Observatory; S. D. Van Dyk,

and K. W. Weiler, Naval Research Laboratory; and N. Panagia, Space Telescope Science Institute, report: “Observations on Mar. 31.07 UT with the Very Large Array in B configuration at 3.6-cm wavelength (8.4 GHz) show no detection to a limit (3σ) of < 0.14 mJy, and at 20-cm wavelength (1.45 GHz) to a limit (3σ) of < 0.16 mJy. Radio monitoring is continuing at these and other wavelengths. For reference purposes, the nucleus of NGC 3031 shows a compact radio source of 65.3 mJy peak, 96.1 mJy integrated, at 1.45 GHz. The position of the nucleus from the 1.45-GHz map is α = 9^{h}51^{m}27^{s}.30, δ = +69^{o}18′07′′.5 (equinox 1950.0; ±0′′.2 in each coordinate).

Weiler, Van Dyk and K. J. Johnston, Naval Research Laboratory, recommend the application of optical astrometric techniques to obtain a highprecision position for SN 1993J during its present optically bright phase. Previous studies of type II supernovae (e.g., Weiler et al. 1986, Ap.J. 301,

790) imply that within a period of weeks to months SN 1993J is very likely to become a significant (> 10 mJy), compact radio source to which radio, including VLBI, astrometric techniques can be applied. Such an astrometric position determination in both wavelength ranges will allow the use of SN 1993J as a radio-optical fiducial point for comparison of the two reference frames as suggested by Weiler et al. (1992, PASP 104, 246).

An Haute Provence spectrum (resolution 0.13 nm) by Y. Andrillat on Mar. 31.87 UT shows the Hα emission to be weaker than on Mar. 30.95 (cf. IAUC 5736). The equivalent width is 0.015 nm with FWHM 145 km/s.

J. C. Wheeler, University of Texas at Austin, reports: “Unreduced optical spectra by E. Barker on the 2.1-m telescope at McDonald on Apr. 1.2 UT are still essentially featureless. D. Lester, N. Gaffney and B. Smith obtained low-resolution K-band spectra with the infrared spectrograph on the 2.7-m telescope that are also featureless; K = 10.4 on Apr. 2.2 UT. Lowresolution J-band spectra gave J = 10.7 on Apr. 2.3. There is now definite evidence for Paschen β, ∼ 4 pixels = 10 000 km/s FWHM. The center of the line is now ∼ 10 percent above continuum with estimated equivalent width 0.002 μm. The early onset of decline in visual bands and the weak features in the spectra are reminiscent of SN 1979C and SN 1980K, both type II linear supernovae. If this is a proper indentification, SN 1993J may become a strong radio source, may retain a strong ultraviolet flux and may show strong He I lines (Swartz, Wheeler and Harkness 1991, Ap.J. 374, 266), but it will decline more rapidly than a type II plateau supernova.”

1993 April 3 Brian G. Marsden